# Distribution-Level Feature Distancing for Machine Unlearning: Towards a Better Trade-off Between Model Utility and Forgetting

### v3

@@ Discussion
A trained model generally shows lower loss values for training data compared to unseen data, which can unintentionally lead to data leakage. Methods like UNSIR (Tarun et al. 2023) and SCRUB (Kurmanji, Triantafillou, and Triantafillou 2023) that maximize loss for data intended to be forgotten may inadvertently increase the loss for forget data beyond that of unseen data, making the model vulnerable to membership inference attacks.
Our findings reveal that even with unlearning, naive error-maximization can still result in information leakage. Specifically, when the number of forget samples is small ($<100$), the loss values for forget data can abnormally increase, exceeding those of unseen data (Figure 6). This issue highlights a risk that has been largely overlooked in prior studies.
Figure 6: The naive instance-level error-maximizing can induce very excessively high loss value for forget data even larger than the unseen data, which might not be desirable.
@@
Our method reveals a trade-off between test accuracy and forgetting score. As we increase the loss for the data intended to be forgotten ($x_{forget}$), the forgetting score improves, but this comes at the cost of test accuracy. This is likely due to correlation collapse, where essential label-related features of the retain data are altered.
Additionally, the effectiveness of unlearning strategies can vary depending on the dataset’s characteristics, scale, and task complexity. This variability suggests the challenges of setting up robust unlearning experiments and the need for adaptive unlearning methods adjusted to different scenarios. Our findings emphasize the importance of carefully balancing model utility and forgetting performance to achieve optimal unlearning results.
@@
Moreover, while the MIA is widely used as a metric to assess forgetting performance, it may not fully capture unlearning effectiveness across all scenarios. In scenarios where the model is exceptionally well-trained, the distinction between forget and unseen data may become minimal, leading to MIA scores that do not adequately reflect the true forgetting performance. This suggests the need for the unlearning community to develop more robust evaluation metrics.
@@ Acknowledgements
This research was supported by Brian Impact, a non-profit organization dedicated to advancing science and technology. We also express our gratitude to MODULABS for their support of our research.

### v4

@@ Discussion
A trained model generally shows lower loss values for training data compared to unseen data, which can lead to data leakage. Methods like UNSIR (Tarun et al. 2023) and SCRUB (Kurmanji, Triantafillou, and Triantafillou 2023) that maximize loss for data intended to be forgotten may inadvertently increase the loss for forget data beyond that of unseen data, making the model vulnerable to membership inference attacks. Our findings reveal that even with unlearning, naive error-maximization can still result in information leakage. Specifically, when the number of forget samples is small ($<100$), the loss values for forget data can abnormally increase, exceeding those of unseen data (Figure 6). This issue highlights a risk that has been overlooked in prior studies.
Figure 6: The naive instance-level error-maximizing can induce excessively high loss value for forget data even larger than the unseen data, which might not be desirable.
@@
Our method reveals a trade-off between test accuracy and forgetting score. As we increase the loss for the data intended to be forgotten ($x_{forget}$), the forgetting score improves, but this comes at the cost of test accuracy. This is likely due to correlation collapse, where essential label-related features of the retain data are altered. Moreover, the effectiveness of unlearning strategies can vary depending on the dataset’s characteristics, scale, and task complexity. This variability suggests the challenges of setting up robust unlearning experiments and the need for adaptive unlearning methods adjusted to different scenarios. Our findings emphasize the importance of carefully balancing model utility and forgetting performance to achieve optimal unlearning results.
@@
Moreover, although MIA is widely used as a metric to assess forgetting performance, it may not fully capture unlearning effectiveness across all scenarios. In scenarios where the model is exceptionally well-trained, the distinction between forget and unseen data may become minimal, leading to MIA scores that do not adequately reflect the true forgetting performance. This suggests the need for the unlearning community to develop more robust evaluation metrics.
@@ Acknowledgements
This research was supported by Brian Impact, a non-profit organization dedicated to advancing science and technology.